OpenStudy (anonymous):

Find an expression for the inverse of the function: f(x)=(2e^x)−1

OpenStudy (anonymous):

Replace x and y. Then solve. x=2e^y-1 x+1=2e^y x/2+1/2=e^y y=ln((1/2)(x+1)) or y=ln(1/2)+ln(x+1)